Major Events

Table of contents
No headers

 

Date Range Event Notes
September 3 - 17, 2012 Laser failure and attempts at realigning the IFS. Installed SIOS laser on Sept 5, and SpectraPhysics on Sept 12. No data.
September 17, 2012 Installed a new InSb filter  
September 20, 2012 First day of fully automated measurements.  
November 2, 2012 Installed lock washer on LASA/LASB detectors. Better laser detector stability/alignment.
January 29, 2013 Installed washers under fixed cube corner locking screws. Better cube corner stability and increased alignment control.
April 2, 2013-May 22, 2013 Refurbushed SIOS laser installed. Laser was not good and significantly increased noise. SpectraPhysics laser reinstalled on May 22, 2013.
August 4, 2014 - August 20, 2014 InSb detector not staying cold. No InSb measurements. InGaAs/Si unaffected.
April 20, 2015 Replaced scanner motor/encoder assembly. Much less noisy measurements. Both forward and reverse can be used independently.
July 10, 2015 Instrument realignment and ghost minimization. See this page for details. Data between July 10-24 have very poor ILS and are manually flagged

 

Tag page